The average expenditure on Valentines Day was expected to be $100.89 USA Today, February 13, 2006). Do male and se ale co s mers drer in the amounts they spen The rage expendere In sample su vey of 58 male consumers was 133, and the average expenditure i sample survey of 35 female consumers was se. Based on past survey, the standard de ation for nale corres assumed to be $35, and the standard deviation for female consumers is assumed to be $10. The z value is 2.576 Round your answers to 2 decimal places a. What is the point estimate of the difference between the population mean expenditure for males and the population mean expenditure for females? b. At 99% ceendence, what is the margin or error? e. Developa99% condence intervenor the dimerence between the two population means 03 63
Answer #1

a)

point estimate of difference =133-69=64

b) margin of error =z*sqrt(\sigma12/n1+\sigma22/n2)=12.61

c)99% CI =64-/+12.61 =51.69 to 76.61